1.The pH sample of human blood was measured to be 7.41 at 25 degrees celsius. Calculate the hydrogen ion concentration, H' and hydroxide ion, OH for the sample.

1
Expert's answer
2022-05-23T17:03:04-0400

Solution:

We can convert between pH and [H+] using the following equation:

pH = −log[H+]

[H+] = 10−pH = 10−7.41 = 3.89×10−8

[H+] = 3.89×10−8 M


For any aqueous solution at 25C:

pH + pOH = 14

pOH = 14 − pH = 14 − 7.41 = 6.59

pOH = 6.59


We can convert between pOH and [OH] using the following equation:

pOH = −log[OH]

[OH] = 10−pOH = 10−6.59 = 2.57×10−7

[OH] = 2.57×10−7 M


Answer:

The hydrogen ion (H+) concentration is 3.89×10−8 M

The hydroxide ion (OH) concentration is 2.57×10−7 M